OpenStudy (anonymous):

What number makes this polynomial a perfect square x^2+15x+k

OpenStudy (anonymous):

\[\frac{225}{4}\]

OpenStudy (anonymous):

seems random i know, but it goes like this. half of 15 is \[\frac{15}{2}\] and \[(\frac{15}{2})^2=\frac{15^2}{2^2}=\frac{225}{4}\]

OpenStudy (anonymous):

- Thats what I thought , im ohkay w/ this stuff , juss not good .

OpenStudy (akshay_budhkar):

you have to express it in the form of a^2+2ab+b^2

OpenStudy (akshay_budhkar):

oh ok...satellite73 explained u.. kk